Diffusion Morning Report 05/30/06

Ted Berg tberg at snf.stanford.edu
Tue May 30 07:53:22 PDT 2006

Good Morning All,
       Well, welcome back everyone. It looks like the end of a long 
weekend. LTO had a broken injector-replaced, Poly is not pumping down
Looking into it, Nitride has DCS flow issues, Epi ran out of HCL it was 
switched over this morning.  Tylan 6 has TCU errors. Applied is here
 working on Tier 0 for the new Epi. Everything else looks OK for now.

Diffusion Mike,Ray, Ted 	  	  	  	  	 
  	Date: 	5/30/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Down
	TCU errors
Tylan 7-TEOS2 	Up 	  	2.7/30 	  	 
Tylan 8-SiGe 	Up 	  	30.17/30 	  	 
Tylan 9-Poly 	Down 	Not Pumping down
Tylan 10-Nitride 	Down 	DCS flow issues 	0/30 	  	 
Tylan 11-BPSG-LTO 	yellow 	poor uniformity 	5.53/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Down 	Testing 	  	  	 
Epi 	yellow 	possible higher leak rate 	  	  	 
Burn Box 	Up 	  	  	  	 
ZYGO 	up 	  	   	  	 
GAAS 22 	UP 	  	  	  	 
GAAS 23 	up 	  	  	  	 
SiH4 for furnaces 	  	A on line 500 psi 	 changeover working well  	  	 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060530/4e43cc22/attachment.html>

More information about the diffusion mailing list